OpenStudy (anonymous):

check for symmetry, y=(x)/(x^2+1) I think thats how you set it up...

OpenStudy (anonymous):

check that this is odd. \[\frac{odd}{even}=odd\] for functions

OpenStudy (anonymous):

Im so confused...

OpenStudy (anonymous):

ok it is symmetric with respect to the origin.

OpenStudy (anonymous):

we can check first with numbers and then with variables.

OpenStudy (anonymous):

let x = 1, \[y=\frac{1}{1^2+1}=\frac{1}{2}\]

OpenStudy (anonymous):

now let x = -1 \[y=\frac{-1}{(-1)^2+1}=\frac{-1}{1+1}=-\frac{1}{2}\]

OpenStudy (anonymous):

this says go right 1, up one half, left one, down one half.

OpenStudy (anonymous):

since we have \[(1,\frac{1}{2})\] and also \[(-1,-\frac{1}{2})\]

OpenStudy (anonymous):

Oh! I see now! Ok, one more... xy^2+10=0, what do you do with the 10?

OpenStudy (anonymous):

are you sure? we should check with variables as well. x = a, get \[y=\frac{a}{a^2+1}\] x = -a get \[y=-\frac{a}{a^2+1}\]

OpenStudy (anonymous):

so symmetric wrt the origin.

OpenStudy (anonymous):

yeah i get it. the x=1 thing helped

OpenStudy (anonymous):

\[xy^2+10=0\] \[xy^2=-10\] \[x=-\frac{10}{y^2}\]

OpenStudy (anonymous):

ahhhh thank you!

OpenStudy (anonymous):

numbers always help

OpenStudy (anonymous):

system is weird ignore last remark.

OpenStudy (anonymous):

Well thank you very much, helps a lot1